### v1

# The Multi-Query Paradox in Zeroth-Order Optimization

Abstract
Zeroth-order (ZO) optimization provides a powerful framework for problems where explicit gradients are unavailable and have to be approximated using only queries to function value. The prevalent single-query approach is simple, but suffers from high estimation variance, motivating a multi-query paradigm to improves estimation accuracy. This, however, creates a critical trade-off: under a fixed budget of queries (i.e. cost), queries per iteration and the total number of optimization iterations are inversely proportional to one another. How to best allocate this budget is a fundamental, under-explored question.
This work systematically resolves this query allocation problem. We analyze two aggregation methods: the de facto simple averaging (ZO-Avg), and a new Projection Alignment method (ZO-Align) we derive from local surrogate minimization. By deriving convergence rates for both methods that make the dependence on the number of queries explicit across strongly convex, convex, non-convex, and stochastic settings, we uncover a stark dichotomy: For ZO-Avg, we prove that using more than one query per iteration is always query-inefficient, rendering the single-query approach optimal. On the contrary, ZO-Align generally performs better with more queries per iteration, resulting in a full-subspace estimation as the optimal approach. Thus, our work clarifies that the multi-query problem boils down to a choice not about an intermediate query size, but between two classic algorithms, a choice dictated entirely by the aggregation method used. These theoretical findings are also consistently validated by extensive experiments.
1 Introduction
Zeroth-order (ZO) optimization provides a powerful framework for problems where explicit gradient information is unavailable or prohibitively expensive to compute (Conn, Scheinberg, and Vicente 2009). Unlike first-order methods that rely on analytical gradients, ZO algorithms guide their search using only function value queries to approximate gradients. Typically, it is achieved by evaluating the objective function along one or more random directions and observing the change in function value (Spall 2002; Mania, Guy, and Recht 2018). This paradigm has recently gained significant attention due to its applicability to a wide array of challenging, high-dimensional problems, including fine-tuning of LLMs (Malladi et al. 2023; Zhang et al. 2024), policy optimization in reinforcement learning (Sun et al. 2020), and hyperparameter tuning (Li et al. 2021).
The predominant approach in modern ZO optimization is the single-query method, where a gradient estimate is constructed using one random direction per optimization step (Nesterov and Spokoiny 2017). While foundational, these estimators are inherently noisy, with an estimation variance that often scales with the problem’s dimension (Nesterov and Spokoiny 2017). A natural extension to mitigate this issue is the multi-query approach, which employs a block of queries at each iteration to construct a more accurate gradient estimate (Duchi et al. 2015). However, this approach introduces a fundamental trade-off. Using multiple queries per iteration inverse proportionally reduces the total number of iteration steps that can be performed under a fixed budget of queries. A systematic investigation of the multi-query paradigm requires addressing two basic questions: (1) What is the most effective method for combining information from multiple query directions into a single, cohesive gradient estimate? (2) For a fixed total query budget, is it more efficient to perform many iterations with noisy, single-query estimates, or fewer iterations with higher-quality, multi-query estimates by consuming more queries per iteration?
This work systematically addresses these questions. First, to resolve the aggregation problem, we propose a new multi-query estimator called Projection Alignment (ZO-Align), from the first principle of local surrogate minimization. While similar estimators have been proposed from the perspective of sparse gradient recovery (Cai et al. 2022; Wang et al. 2018), our optimization-centric derivation provides a more fundamental justification for the method, independent of any sparsity assumptions. For a comprehensive comparison, we also analyze the canonical strategy of simply averaging independent estimates (ZO-Avg) (Duchi et al. 2015). Our initial analysis of the estimators’ properties reveals that ZO-Align achieves a significantly lower mean squared error, motivating a deeper investigation into its performance within an optimization context.
Second, to resolve the query allocation problem, we derive the convergence rates for algorithms using both ZO-Align and ZO-Avg, making the dependence on the number of queries used per iteration explicit. We then leverage these rates to formally pose and solve an optimization problem: for a fixed total query budget, find the best allocation of queries per iteration that minimizes the final optimization error. To ensure the generality and robustness of our conclusions, this entire theoretical framework is systematically applied across a wide spectrum of problem structures, including strongly convex, convex, non-convex, and stochastic settings.
Our analysis reveals a stark dichotomy where the optimal allocation strategy is critically dependent on the chosen aggregation method. For ZO-Avg, we prove that it cannot effectively utilize information from multiple queries; any attempt to use more than one query per iteration is detrimental, and the optimal strategy is always the single-query approach. In direct contrast, the ZO-Align estimator is capable of efficiently harnessing multi-query information. For ZO-Align, we prove that performance improves with the number of queries, making the optimal strategy a full-subspace estimation where the number of queries matches the problem’s dimension.
To complement our theoretical analysis, we conducted extensive experiments across a suite of optimization problems. The empirical results consistently validate our theoretical findings, demonstrating the practical implications of the query allocation dichotomy. Our work clarifies that the choice is not about finding an optimal intermediate query size, but rather a commitment to an aggregation method that dictates whether the optimal algorithm is a sequential single-query method or a full-subspace estimation.
The primary contributions of this work are as follows:
  • We propose the ZO-Align estimator based on local surrogate minimization, providing a clear optimization-centric perspective that complements existing geometric interpretations.
  • We deliver the first comprehensive convergence analysis that directly compares ZO-Avg and ZO-Align across strongly convex, convex, non-convex, and stochastic settings, revealing a fundamental and consistent dichotomy in their performance.
  • We provide rigorous theoretical proofs that under a fixed query budget, the optimal query allocation for ZO-Avg is always single-query, whereas the optimal allocation for ZO-Align is to use the maximum possible number of queries, resolving the query allocation question for these two classes of estimators.
  • We empirically verify our theoretical findings, demonstrating that ZO-Align with a large query size significantly outperforms both ZO-Avg and single-query baselines across various optimization settings.
2 Related Work
ZO gradient estimation.
Zeroth-order optimization approximates gradients using finite-difference techniques, a concept explored through various random perturbation strategies. For instance, theoretical convergence has been demonstrated using Gaussian random noise perturbations (Nesterov and Spokoiny 2017), while other approaches have utilized uniform sampling (Flaxman, Kalai, and McMahan 2004) and coordinate-wise perturbations (Lian et al. 2016). A distinct line of research focuses on problems where the gradient is assumed to be sparse, framing the estimation task as one of compressed sensing; here, techniques like LASSO (Wang et al. 2018) or other forms of regularized regression (Choromanski et al. 2020) enable the reconstruction of a sparse gradient from a limited number of function queries (Cai et al. 2022). More recently, methods based on linear interpolation have been investigated as a potentially more accurate alternative (Wang et al. 2024). Research suggests that linear interpolation can yield better gradient estimates than Gaussian smoothing (Berahas et al. 2019) and may require significantly fewer samples to do so (Berahas et al. 2022).
Query-efficient ZO optimizers.
A key challenge in ZOO is its high query complexity, motivating research into more efficient methods. Some methods build more accurate local models to estimate update direction by incorporating second-order Hessian information (Ye et al. 2025), or by leveraging problem geometry via Riemannian optimization (He et al. 2024). Other approaches optimize the querying process itself, such as using compressed sensing to achieve double-logarithmic query complexity for sparse problems (Qiu and Tong 2024), adaptively reusing old queries to reduce variance (Xiao, Ling, and Chen 2023), or learning a sampling policy with reinforcement learning (Zhai et al. 2024). However, despite these advances, these approaches are all centered on improving the quality of the estimate at each individual step. In contrast, the fundamental problem of how to optimally allocate a fixed query budget to maximize its utility remains a critical yet under-explored question.
3 ZO Estimators: From Single-Query to Multi-Query
This section introduces the core ZO gradient estimators used in our analysis. We begin with the standard single-query estimator before presenting two multi-query approaches: the intuitive Naive Averaging (ZO-Avg) and the principled Projection Alignment (ZO-Align), which we derive from local surrogate minimization. Then we analyze their Mean Squared Error (MSE) to establish a fundamental difference in their accuracy, which underpins the main convergence results in Section 4.
3.1 Preliminaries and the Single-Query Estimator
We consider the unconstrained optimization problem $\min_{x\in\mathbb{R}^{d}}f(x)$ where direct access to the gradient $\nabla f(x)$ is unavailable.. Throughout this work, we assume the objective function $f$ is differentiable and $L$-smooth:
$\|\nabla f(x)-\nabla f(y)\|\leq L\|x-y\|,\quad\forall x,y\in\mathbb{R}^{d}.$
ZO algorithms navigate this problem by iteratively updating a parameter vector $x_{t}$ using a gradient estimate $\hat{g}(x_{t})$ constructed from function value queries: $x_{t+1}=x_{t}-\eta_{t}\hat{g}(x_{t})$.
The foundational building block for most ZO methods is the single-query estimator. It approximates the gradient by evaluating the function along a single random direction $u\in\mathbb{R}^{d}$. A common form is given by:
$\hat{g}_{u}(x)=\frac{f(x+\mu u)-f(x)}{\mu}u,$
where $\mu>0$ is a small smoothing parameter. This form is derived from the first-order Taylor expansion of $f(x)$ around $x$:
$f(x+\mu u)=f(x)+\mu u^{T}\nabla f(x)+O(\mu^{2}).$
For a sufficiently small $\mu$, the estimator approximates the projection of the true gradient onto the direction $u$: $\hat{g}_{u}(x)\approx(u^{T}\nabla f(x))u$. While fundamental, this single-direction estimate is inherently noisy, limiting its effectiveness. A natural extension is to probe the function along multiple directions to construct a more accurate gradient estimate.
3.2 Simple Averaging Estimator (ZO-Avg)
The most direct method for aggregating information from a block of $q$ queries is to compute $q$ independent single-query estimates and average them. Let $U=[u_{1},\dots,u_{q}]\in\mathbb{R}^{d\times q}$ be a matrix whose columns are random direction vectors drawn independently from $\mathcal{N}(0,I)$. We let $q\leq d$ and assume the columns of $U$ are linearly independent. The averaging estimator, which we term ZO-Avg, is defined as:
$\hat{g}_{\text{AVG}}(x)=\frac{1}{q}\sum_{i=1}^{q}\frac{f(x+\mu u_{i})-f(x)}{\mu}u_{i}.$
Our subsequent theoretical analysis focuses on the idealized estimators where the bias from the finite difference is negligible (i.e., as $\mu\to 0$). This allows us to isolate the error stemming from the geometric properties of the estimators. In this idealized form, the estimator approximates:
$\hat{g}_{\text{AVG}}(x)\approx\left(\frac{1}{q}\sum_{i=1}^{q}u_{i}u_{i}^{T}\right)\nabla f(x)=\frac{1}{q}UU^{T}\nabla f(x).$
This estimator is appealing due to its simplicity and straightforward parallel implementation. A key property of ZO-Avg is that it provides an unbiased estimate of the true gradient.
Proposition 3.1.
Let the columns of $U$ be i.i.d. samples from $\mathcal{N}(0,I)$. The ZO-Avg estimator is unbiased, i.e., $\mathbb{E}[\hat{g}_{AVG}(x)]=\nabla f(x)$.
Proof.
See Appendix A.1. ∎
While unbiasedness is a desirable property, it does not guarantee efficiency. As we will show, the high variance of this estimator plays a critical role in its convergence behavior.
3.3 Projection Alignment Method (ZO-Align)
Instead of simply averaging, we adopt a more principled, model-based perspective. The $q$ queries provide a snapshot of the function’s local geometry within the column space of $U$. We seek the best update step of the form Uy-Uy for some coefficient vector $y\in\mathbb{R}^{q}$ by minimizing a local surrogate of the objective function. Given that $f$ is $L$-smooth, we can form a local quadratic upper bound around the point $x$:
$f(x-Uy)\leq f(x)-\nabla f(x)^{T}(Uy)+\frac{L}{2}\|Uy\|^{2}.$
Minimizing this quadratic upper bound with respect to $y$ is a principled way to find the best local step within the column space. This yields the optimal coefficient vector:
$y^{*}=\frac{1}{L}(U^{T}U)^{-1}U^{T}\nabla f(x).$
The optimal update direction is therefore proportional to $Uy^{*}=U(U^{T}U)^{-1}U^{T}\nabla f(x)$. This motivates our second estimator, which replaces the true directional derivatives with their finite-difference approximations:
$\hat{g}_{{\text{ALG}}}(x)=U(U^{T}U)^{-1}\begin{bmatrix}\frac{f(x+\mu u_{1})-f(x)}{\mu}\\ \vdots\\ \frac{f(x+\mu u_{q})-f(x)}{\mu}\end{bmatrix}$.
The name “Projection Alignment” comes from the powerful geometric property of this estimator. The matrix $P_{U}=U(U^{T}U)^{-1}U^{T}$ is the orthogonal projection operator onto the column space of $U$. Therefore, assuming $\mu\to 0$, $\hat{g}_{\text{ALG}}$ is the orthogonal projection of the true gradient onto the query subspace. This means that, by construction, its own projection onto each query direction $u_{i}$ is identical to the estimated projection of the true gradient: $u_{i}^{T}\hat{g}_{\text{ALG}}\approx u_{i}^{T}\nabla f(x)$.
This reveals a crucial difference from ZO-Avg. While ZO-Align is constructed to be geometrically consistent within the sampled subspace, ZO-Avg is not. The projection of the ZO-Avg estimate onto a query direction $u_{i}$ does not generally match the true directional derivative, leading to a less accurate representation of the local geometry.
Unlike ZO-Avg, ZO-Align is a biased estimator.
Proposition 3.2.
Let the columns of $U$ be i.i.d. samples from $\mathcal{N}(0,I)$. The ZO-Align estimator has an expected value of $\mathbb{E}[\hat{g}_{ALG}(x)]=\frac{q}{d}\nabla f(x)$.
Proof.
See Appendix A.1. ∎
3.4 Accuracy Analysis: A Tale of Two Estimators
The two estimators present a classic statistical trade-off. ZO-Avg is unbiased but suffers from high variance. In contrast, ZO-Align accepts a well-structured bias in exchange for a dramatic reduction in estimation error. To formalize this comparison, we analyze the Mean Squared Error (MSE) of each estimator, defined as $\mathbb{E}[\|\hat{g}(x)-\nabla f(x)\|^{2}]$, which captures the total error from both bias and variance.
Proposition 3.3.
Let $g=\nabla f(x)$ and assume the query directions are i.i.d. samples from $\mathcal{N}(0,I)$. The Mean Squared Error of the idealized estimators are:
  1. 1.
    For ZO-Avg: $\quad\text{MSE}(\hat{g}_{\text{AVG}})=\frac{d+1}{q}\|g\|^{2}$.
  2. 2.
    For ZO-Align: $\quad\text{MSE}(\hat{g}_{\text{ALG}})=\frac{d-q}{d}\|g\|^{2}$.
Proof.
See Appendix A.1. ∎
For any $q>1$, it is clear that $\frac{d-q}{d}\ll\frac{d+1}{q}$, indicating that ZO-Align offers a significantly lower estimation error. This fundamental difference in estimator quality is the primary driver of the performance gap we will establish in the following sections. It highlights that the principled, model-based approach of ZO-Align is far more effective at aggregating information from multiple queries than the naive averaging of ZO-Avg.
4 Query Allocation Dichotomy
This section presents our main theoretical results, which establish a fundamental dichotomy in how multi-query ZO estimators should be utilized. Upon several standard settings, we analyze the convergence behavior of optimization algorithms using the ZO-Avg and ZO-Align estimators under a fixed total query budget. We first consider the deterministic setting, where we have exact function evaluations, for three classes of functions: $\mu$-strongly convex, convex, and non-convex. We then extend our analysis to the stochastic setting for convex functions, where function evaluations are corrupted by noise. Across all these settings, we show that the simple averaging of ZO-Avg is an inefficient mechanism for aggregating query information, making a fully sequential single-query method optimal. In direct contrast, the principled construction of ZO-Align effectively harnesses multi-query information, making the optimal strategy a full-subspace estimation.
4.1 Query Allocation Problem
Suppose a fixed total query budget, $K$, an algorithm performs a series of updates over $T$ iterations, and at each iteration $t$, it uses a set of $q_{t}$ function queries as stated in Section 3.2. The total number of queries is therefore constrained by $\sum_{t=0}^{T-1}q_{t}\leq K$. Our central goal is to determine the optimal query allocation strategy—the sequence of queries per iteration, $\{q_{t}\}_{t=0}^{T-1}$—that minimizes the final optimization error.
4.2 Optimal Query Allocation
We begin by analyzing the case where function evaluations are deterministic. The update rule at iteration $t$ is $x_{t+1}=x_{t}-\eta_{t}\hat{g}(x_{t})$, where $\hat{g}$ is either the ZO-Avg or ZO-Align estimator constructed with a set of $q_{t}$ query directions, as illustrated in Section 3.
Strongly Convex Case
A function is defined as $\mu$-strongly convex if there exists a constant $\mu$ such that for all $x$ and $y$ in its domain:
$f(y)\geq f(x)+\nabla f(x)^{T}(y-x)+\frac{\mu}{2}\|y-x\|^{2}.$
When the objective function $f(x)$ is both $L$-smooth and $\mu$-strongly convex, we can achieve a linear convergence rate (Boyd and Vandenberghe 2004). However, the rate’s dependence on the number of queries per iteration, $q_{t}$, reveals the core dichotomy of our work. We first analyze the ZO-Avg estimator.
Theorem 4.1.
Let $f$ be an $L$-smooth and $\mu$-strongly convex function. The ZO-Avg algorithm with the optimal step size $\eta_{t}=\frac{q_{t}}{L(q_{t}+d+1)}$ achieves the following convergence guarantee:
$\begin{split}&\quad\mathbb{E}[f(x_{T})-f(x^{*})]\\ \leq&\left(\prod_{t=0}^{T-1}\left(1-\frac{\mu q_{t}}{L(q_{t}+d+1)}\right)\right)(f(x_{0})-f(x^{*})),\end{split}$
where $x^{*}$ represent the optimum. Under a fixed total query budget $K$, this convergence rate is optimized by choosing $q_{t}=1$ for all $t=0,\dots,K-1$, resulting in $T=K$ iterations and a rate of:
$\mathbb{E}[f(x_{K})-f(x^{*})]\leq\left(1-\frac{\mu}{L(d+2)}\right)^{K}(f(x_{0})-f(x^{*}))$
Proof.
See Appendix A.1. ∎
The result in Theorem 4.1 is deeply counter-intuitive. It suggests that when using the simple averaging estimator, allocating more than one query per iteration is detrimental to the convergence rate. The core reason lies in the trade-off between variance reduction and the cost of queries. While increasing $q_{t}$ does reduce the variance of the gradient estimate, the improvement scales as $1/q_{t}$ as shown in Section 3.4. However, the cost increases linearly with $q_{t}$. The analysis shows that the marginal benefit of reducing variance is not worth the cost of the additional queries. It is always better to spend the query budget on taking more, albeit noisier, steps. This exposes a fundamental flaw in the simple average strategy: it is an inefficient use of a limited query budget.
In stark contrast, the principled structure of the ZO-Align estimator leads to the opposite conclusion.
Theorem 4.2.
Let $f$ be an $L$-smooth and $\mu$-strongly convex function. The ZO-Align algorithm with the optimal step size $\eta_{t}=1/L$ achieves the following convergence guarantee:
$\mathbb{E}[f(x_{T})-f(x^{*})]\leq\left(\prod_{t=0}^{T-1}\left(1-\frac{\mu q_{t}}{Ld}\right)\right)(f(x_{0})-f(x^{*})),$
where $x^{*}$ represent the optimum. Under a fixed total query budget $K$, this convergence rate is optimized by using as few iterations as possible with the maximum number of queries, i.e., $q_{t}=d$ for $T=K/d$ iterations. This yields the rate:
$\mathbb{E}[f(x_{K/d})-f(x^{*})]\leq\left(1-\frac{\mu}{L}\right)^{K/d}(f(x_{0})-f(x^{*})).$
Proof.
See Appendix A.1. ∎
Theorem 4.2 presents a picture opposite to that of ZO-Avg. For ZO-Align, the optimal strategy is to use the maximum number of queries per step. The intuition is that ZO-Align does not simply average away noise; it constructs a principled, low-error projection of the gradient onto the query subspace. As shown in Section 3.4, the quality of this projection improves substantially as the number of query directions increases. Each additional query provides a significant marginal benefit by better aligning the search direction with the true gradient. It is therefore more effective to spend the budget forming a single high-quality, confident step direction than to take many small, uncertain steps. This result highlights the power of a model-based approach: by aggregating information correctly, investing more queries per step becomes a powerful tool for accelerating convergence.
Convex Case
We then analyze the case where the objective function $f(x)$ is convex and $L$-smooth, but not necessarily strongly convex. We demonstrate that the optimal query allocation strategy is fundamentally different for the two estimators, reinforcing the core dichotomy of our work.
Theorem 4.3.
Let $f$ be an $L$-smooth and convex function. The ZO-Avg algorithm with step size $\eta_{t}=\frac{q_{t}}{L(q_{t}+d+1)}$ achieves the following convergence guarantee after $T$ iterations:
$\begin{split}E[f(x_{T})-f(x_{*})]\leq\frac{L(\|x_{0}-x_{*}\|^{2}+\frac{2}{L}(f(x_{0})-f(x_{*})))}{\sum_{t=0}^{T-1}\frac{2q_{t}}{q_{t}+d+1}}.\end{split}$
Under a fixed total query budget $K$, this convergence rate is optimized by choosing $q_{t}=1$ for all t=0,,T1t=0,...,T-1, resulting in $T=K$ iterations and a rate of:
$\begin{split}&\quad E[f(x_{K})-f(x_{*})]\\ &\leq\frac{L(d+2)(\|x_{0}-x_{*}\|^{2}+\frac{2}{L}(f(x_{0})-f(x_{*})))}{2K}.\end{split}$
Proof.
See Appendix A.1. ∎
Theorem 4.3 reinforces the conclusion from the strongly convex case: for the ZO-Avg estimator, using multiple queries per step is an inefficient use of the query budget. Next, we show that for the ZO-Align estimator, the conclusion remains the opposite.
Theorem 4.4.
Let $f$ be an $L$-smooth and convex function. The ZO-Align algorithm with step size $\eta_{t}=1/L$ achieves the following rate:
$\begin{split}&\quad E[f(x_{T})-f(x_{*})]\\ &\leq\frac{d}{2\sum_{t=0}^{T-1}q_{t}}(L\|x_{0}-x_{*}\|^{2}+2(f(x_{0})-f(x_{*})))\end{split}$
Under a fixed total query budget $K$, ZO-Align yields the rate:
$E[f(x_{T})-f(x_{*})]\leq\frac{d}{2K}(L\|x_{0}-x_{*}\|^{2}+2(f(x_{0})-f(x_{*})))$
under any allocation $\{q_{t}\}$ such that $\sum_{t=0}^{T-1}q_{t}=K$.
Proof.
See Appendix A.1. ∎
Comparing the rates for ZO-Avg and ZO-Align, we see a factor of $d+2$ versus $d$. More importantly, the ZO-Align rate depends only on the total number of queries $K$, not on how they are allocated across iterations. This key distinction stems directly from the mathematical properties of the estimators. The per-iteration progress for ZO-Align is linear in the number of queries. However for ZO-Avg, it is a nonlinear concave function.
Non-Convex Case
Then we proceed to non-convex setting, where the objective is only assumed to be $L$-smooth. In this scenario, the goal is typically to find a first-order stationary point (Nocedal and Wright 2006), i.e., a point $x$ where $\|\nabla f(x)\|^{2}\leq\epsilon$. The convergence is measured by the number of queries required to achieve this goal. The analysis reveals the same fundamental dichotomy between the two estimators.
Theorem 4.5.
Let $f$ be an $L$-smooth function. The ZO-Avg algorithm with step size $\eta_{t}=\frac{q_{t}}{L(q_{t}+d+1)}$ ensures that:
$\min_{t=0,...,T-1}\mathbb{E}[\|\nabla f(x_{t})\|^{2}]\leq\frac{2L(f(x_{0})-f^{*})}{\sum_{t=0}^{T-1}\frac{q_{t}}{q_{t}+d+1}},$
where $f^{*}$ is the minimum value of $f$. Under a fixed total query budget $K$, this rate is optimized by choosing $q_{t}=1$ for all t=0,,T1t=0,...,T-1, which gives $T=K$ iterations and yields:
$\min_{t=0,...,K-1}\mathbb{E}[\|\nabla f(x_{t})\|^{2}]\leq\frac{2L(d+2)(f(x_{0})-f^{*})}{K}$
Proof.
See Appendix A.1. ∎
Theorem 4.6.
Let $f$ be an $L$-smooth function. The ZO-Align algorithm with step size $\eta_{t}=1/L$ ensures that:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(x_{t})\|^{2}]\leq\frac{2Ld(f(x_{0})-f^{*})}{\sum_{t=0}^{T-1}q_{t}}$
Under a fixed total query budget $K$, any allocation $\{q_{t}\}$ that exhausts the budget, $\sum_{t=0}^{T-1}q_{t}=K$, yields the rate:
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla f(x_{t})\|^{2}]\leq\frac{2Ld(f(x_{0})-f^{*})}{K}$
Proof.
See Appendix A.1. ∎
The conclusions for the non-convex setting mirror those from the convex case, reinforcing the same central dichotomy. While the same fundamental difference in optimal allocation holds, the performance gap between the two methods is less pronounced here than in the strongly convex setting, boiling down to a constant factor difference rather than an exponential one.
Stochastic Case
We now extend our analysis to the stochastic setting, a standard framework in modern machine learning where the objective is an expectation: $f(x)=\mathbb{E}_{\xi}[F(x,\xi)]$. At each iteration $t$, the true function $f(x_{t})$ is inaccessible. Instead, our zeroth-order estimators must be constructed from a single stochastic realization $F(x_{t},\xi_{t})$ using a random sample $\xi_{t}$. Our zeroth-order estimators are then constructed using this stochastic function. We show that the fundamental dichotomy between ZO-Avg and ZO-Align persists even in the presence of this new source of noise.
For this analysis, we adopt standard assumptions: $f(x)$ is convex, each component function $F(x,\xi)$ is L-smooth, and the stochastic gradients have bounded variance at the optimum, $x^{*}=\arg\min_{x}f(x)$.
Assumption 4.7.
The variance of the stochastic gradient at the optimum is bounded, i.e.,
$\mathbb{E}_{\xi}[\|\nabla F(x^{*},\xi)-\nabla f(x^{*})\|^{2}]=\mathbb{E}_{\xi}[\|\nabla F(x^{*},\xi)\|^{2}]\leq\sigma^{2}.$
This leads to the following standard lemma which bounds the expected squared norm of the stochastic gradient.
Lemma 4.8.
For a convex and L-smooth function $f(x)$ satisfying Assumption 4.7, we have $\mathbb{E}_{\xi}[\|\nabla F(x,\xi)\|^{2}]\leq 4L(f(x)-f(x^{*}))+2\sigma^{2}$.
Proof.
See Appendix A.1. ∎
This lemma provides the key tool for controlling the error introduced by the stochastic oracle, which we now use to derive the convergence guarantee for two estimators.
Theorem 4.9.
Let $f$ be an L-smooth and convex function satisfying Assumption 4.7. The ZO-Avg algorithm using a diminishing step size $\eta_{t}=\eta_{0}/\sqrt{t+1}$, with $\eta_{t}\leq\frac{1}{4L}$, achieves the following convergence guarantee after $T$ iterations:
$\mathbb{E}[f(\bar{x}_{T})-f(x^{*})]\leq\frac{\|x_{0}-x^{*}\|^{2}+2\eta_{0}^{2}\sigma^{2}\sum_{t=0}^{T-1}\frac{q_{t}+d+1}{q_{t}(t+1)}}{\eta_{0}\sum_{t=0}^{T-1}\frac{d+1-q_{t}}{q_{t}\sqrt{t+1}}}$
where $\bar{x}_{T}$ is the weighted average of the iterates. Under a fixed total query budget $K$, this convergence rate is optimized by choosing $q_{t}=1$ for all $t=0,\dots,K-1$, resulting in $T=K$ iterations.
Proof.
See Appendix A.1. ∎
Theorem 4.10.
Let $f$ be an L-smooth and convex function satisfying Assumption 4.7. The ZO-Align algorithm using a diminishing step size $\eta_{t}=\eta_{0}/\sqrt{t+1}$, with $\eta_{t}\leq\frac{1}{4L}$, achieves the following convergence guarantee after $T$ iterations:
$\mathbb{E}[f(\bar{x}_{T})-f(x^{*})]\leq\frac{d\|x_{0}-x^{*}\|^{2}+2\eta_{0}^{2}\sigma^{2}\sum_{t=0}^{T-1}\frac{q_{t}}{t+1}}{\eta_{0}\sum_{t=0}^{T-1}\frac{q_{t}}{\sqrt{t+1}}}$
where $\bar{x}_{T}$ is the weighted average of the iterates. Under a fixed total query budget $K$, this convergence rate is optimized by choosing $q_{t}=d$ for $T=K/d$ iterations.
Proof.
See Appendix A.1. ∎
4.3 The Multi-Query Paradox
Our theoretical analysis across strongly convex, convex, non-convex, and stochastic functions reveals a consistent and powerful dichotomy. The effective use of multiple queries per step is critically dependent on the method used to aggregate the information from those queries. This suggests that the common pursuit of an optimal intermediate block size $q$ where ($1<q<d$) may be misguided. Instead, the landscape of multi-query zeroth-order optimization is defined by two fundamentally different, opposing poles of operation.
A deeper insight emerges when comparing the performance gap between the two estimators across the different function classes. The superiority of ZO-Align is most pronounced in the strongly convex setting and diminishes as the problem structure weakens to convex and non-convex. This is not a coincidence, but rather a reflection of how the problem’s geometry dictates the value of a high-quality gradient estimate. For strongly convex functions, the analysis relies on a multiplicative decrease in error at each step, where the contraction factor is directly improved by a better gradient estimate. The benefit of ZO-Align’s superior update direction compounds exponentially, leading to a vastly different convergence rate.
In contrast, the analyses for convex and non-convex functions rely on weaker potential-based arguments that result in a telescoping sum over the iterations. This framework captures cumulative, additive progress rather than a compounding, multiplicative effect. Within this additive framework, the step-by-step superiority of ZO-Align is averaged out over the entire optimization run. Its advantage manifests only as a slightly better constant factor in the final rate—a factor that becomes marginal for large $d$. The lack of strong geometric structure prevents the algorithm from fully capitalizing on the superior search direction in a compounding manner.
Our analysis of the stochastic convex case further reinforces the central thesis established in the deterministic setting: leveraging parallel queries in ZO methods collapses to a choice between two canonical algorithms. The introduction of noise from the stochastic oracle does not create a new middle ground; it simply reinforces the existing poles.
This trade-off between statistical efficiency and the cost of queries pushes the optimal strategy for each estimator to an extreme. For the ZO-Avg estimator, our theory proves that the optimal strategy is always to set $q_{t}=1$, which reduces the algorithm to a simple, sequential single-query approach. Conversely, for the statistically superior ZO-Align estimator, the optimal strategy is to use the maximum number of queries, setting $q_{t}=d$. In this important special case, the projection matrix $P_{u}=U(U^{T}U)^{-1}U^{T}$ becomes the identity matrix $I$ for any full rank $U$. The ZO-Align estimator thereby converges to the true gradient, and the method transforms into a classical full-subspace finite-difference method.
Therefore, our analysis suggests that the question of how to leverage multiple queries in ZO methods does not lead to a novel intermediate-query algorithm. Rather, it collapses the problem to a choice between two well-established canonical algorithms: a sequential single-query method (the optimal realization of ZO-Avg) and a full-subspace finite-difference method (the optimal realization of ZO-Align). Our work clarifies that these two poles are the theoretically-grounded optimal choices, and the path taken depends entirely on the chosen aggregation strategy.
5 Experiments
Figure 1: Objective with query used: Strongly convex case.
Figure 2: Objective with query used: Convex case.
Figure 3: Objective with query used: Nonconvex case.
Figure 4: Objective with query used: Stochastic case.
To complement our theoretical analysis, we conduct a series of experiments designed to investigate the practical performance implications of the ZO-Avg and ZO-Align estimators across several classic optimization problems. The results provide a clear, practical demonstration of the query allocation dichotomy that forms the central thesis of our work.
5.1 Experimental Setup
Objective Functions. We evaluate the algorithms on four classic objective functions representing strongly convex, convex, non-convex, and stochastic convex problems, with the dimension fixed at $d=1000$.
  • Strongly Convex: A quadratic function, $f(x)=\frac{1}{2}x^{T}Ax+b^{T}x$.
  • Convex: A logistic regression function on synthetic data, $f(x)=\frac{1}{m}\sum_{i=1}^{m}\log(1+\exp(-y_{i}a_{i}^{T}x))$.
  • Non-Convex: The Rosenbrock function (Rosenbrock 1960), $f(x)=\sum_{i=1}^{d-1}[100(x_{i+1}-x_{i}^{2})^{2}+(1-x_{i})^{2}]$.
  • Stochastic Convex: A stochastic logistic regression objective, $f(x)=\mathbb{E}_{(\mathbf{w},y)\sim\mathcal{D}}[\log(1+\exp(-y\mathbf{w}^{T}x))]+\frac{\rho}{2}\|x\|^{2}$. The function is computed using mini-batches of data sampled at each iteration.
Evaluation Metric. We evaluate ZO-Avg and ZO-Align with various query sizes per iteration (q), plotting the objective value against the number of queries used. Each curve represents the average of 10 independent runs for reliability. We test under two total query budgets: a large budget of $K=20,000$ and a constrained budget of $K=500$. The latter case, where the total budget is less than the problem dimension ($K<d$), is crucial for examining the performance dichotomy in a resource-constrained setting where full-subspace estimation is impossible. Other detailed settings are provided in Appendix A.2.
5.2 Empirical Results and Discussion
In this section, we present the empirical results of our experiments, which validate the theoretical dichotomy established in Section 4.3. The following figures illustrate the performance of the ZO-Avg and ZO-Align estimators on various objective functions. For these figures, the horizontal axis represents the number of queries used, while the vertical axis shows the objective value achieved. With the same number of query used, increasing the number of queries per iteration (q) inversely reduces the total number of iterations. To highlight the focus of our research, we use the total query number used rather than the number of iterations to measure the optimization cost, as this more directly reflects the computational efficiency of zeroth order methods.
As shown in Figure 2 for the strongly convex case, ZO-Align with a large query block size is dramatically superior, while ZO-Avg’s performance degrades as $q$ increases. This holds for both the large ($K=20,000$) and small ($K=500$) budgets. The small-budget scenario is particularly revealing: even when the total budget is too small to afford a single full-subspace estimate ($K<d$), ZO-Align with the largest feasible $q$ still outperforms other strategies, highlighting the robust benefit of its aggregation method. For the convex and non-convex problems, Figures 2 and 4 show that this same fundamental pattern persists. However, the performance landscape is more nuanced, as the improvement of the best ZO-Align strategy over the baselines is more modest than in the strongly convex case. This aligns with our theory, which posits that without the compounding effect of strong convexity, the benefit manifests as an improved constant factor rather than an exponential separation.
The stochastic convex setting, shown in Figure 4, further confirms the robustness of our central thesis, as the core dichotomy between the estimators persists even in the presence of noise from the objective function itself. ZO-Align with large $q$ greatly outperform other competitors.
In summary, the experiments confirm that the multi-query optimization challenge collapses to a choice between two fundamentally different algorithms, with the aggregation method as the critical factor. ZO-Avg is an inefficient use of a query budget, making a sequential single-query method its optimal implementation. In contrast, ZO-Align effectively leverages multiple queries, with its performance systematically improving as the query block size increases.
6 Conclusions
This work addresses two core questions in zeroth-order optimization: how to aggregate multi-query information and how to allocate a query budget. We analyze two strategies: simple averaging (ZO-Avg) and a principled Projection Alignment (ZO-Align) method derived from local surrogate minimization. Our analysis, validated across numerous theoretical settings and experiments, reveals a stark dichotomy. We prove the optimal strategy for ZO-Avg is always a sequential random search, while ZO-Align’s performance consistently improves as more queries are used, culminating in a full-subspace estimation. This resolves the multi-query paradox, showing the choice is not about an intermediate query size, but about committing to an aggregation method that determines one of these two canonical algorithms.
Appendix A Appendix
A.1 Proofs
Proof of Proposition 3.1
Proof.
Taking the expectation over the random directions $u_{i}$, we have $\mathbb{E}[u_{i}u_{i}^{T}]=I$. By linearity of expectation, $\mathbb{E}[\frac{1}{q}UU^{T}]=\frac{1}{q}\sum_{i=1}^{q}\mathbb{E}[u_{i}u_{i}^{T}]=\frac{1}{q}\sum_{i=1}^{q}I=I$. Thus, $\mathbb{E}[\hat{g}_{AVG}(x)]=\mathbb{E}[\frac{1}{q}UU^{T}]\nabla f(x)=\nabla f(x)$. ∎
Proof of Proposition 3.2
Proof.
The expectation of a projection matrix onto a random $q$-dimensional subspace in $\mathbb{R}^{d}$ is known to be $\mathbb{E}[P_{U}]=\frac{q}{d}I$. Therefore, $\mathbb{E}[\hat{g}_{LS}(x)]=\mathbb{E}[P_{U}]\nabla f(x)=\frac{q}{d}\nabla f(x)$. ∎∎
Proof of Proposition 3.3
Proof.
Analysis of ZO-Avg. Since ZO-Avg is unbiased, its MSE is the trace of its covariance matrix. Let $g=\nabla f(x)$ and $\hat{g}_{i}=(g^{T}u_{i})u_{i}$. The ZO-Avg estimator is $\hat{g}_{AVG}=\frac{1}{q}\sum_{i=1}^{q}\hat{g}_{i}$. As the directions $u_{i}\sim\mathcal{N}(0,I)$ are i.i.d., the terms $\hat{g}_{i}$ are also i.i.d. random vectors. The variance of the average is thus:
$\text{Var}(\hat{g}_{AVG})=\frac{1}{q^{2}}\sum_{i=1}^{q}\text{Var}(\hat{g}_{i})=\frac{1}{q}\text{Var}(\hat{g}_{1}).$
To compute $\text{Var}(\hat{g}_{1})$, we need the first two moments of $\hat{g}_{1}$. We have $\mathbb{E}[\hat{g}_{1}]=\mathbb{E}[u_{1}u_{1}^{T}]g=Ig=g$. For the second moment, we use the known result for a standard Gaussian vector $u\sim\mathcal{N}(0,I)$ and a fixed vector $v$: $\mathbb{E}[(v^{T}u)^{2}uu^{T}]=2vv^{T}+\|v\|^{2}I$. Thus:
$\mathbb{E}[\hat{g}_{1}\hat{g}_{1}^{T}]=\mathbb{E}[(g^{T}u_{1})^{2}u_{1}u_{1}^{T}]=2gg^{T}+\|g\|^{2}I.$
The covariance matrix of a single estimator is:
$\text{Var}(\hat{g}_{1})$ $=\mathbb{E}[\hat{g}_{1}\hat{g}_{1}^{T}]-\mathbb{E}[\hat{g}_{1}]\mathbb{E}[\hat{g}_{1}]^{T}$
$=(2gg^{T}+\|g\|^{2}I)-gg^{T}=gg^{T}+\|g\|^{2}I.$
The covariance of the ZO-Avg estimator is $\text{Var}(\hat{g}_{AVG})=\frac{1}{q}(gg^{T}+\|g\|^{2}I)$. The MSE is the trace of this matrix:
$\text{MSE}(\hat{g}_{AVG})$ $=\text{tr}(\text{Var}(\hat{g}_{AVG}))=\frac{1}{q}\text{tr}(gg^{T}+\|g\|^{2}I)$
$=\frac{1}{q}(\text{tr}(gg^{T})+\|g\|^{2}\text{tr}(I))=\frac{1}{q}(\|g\|^{2}+d\|g\|^{2})=\frac{d+1}{q}\|g\|^{2}.$
Analysis of ZO-Align. The ZO-Align estimator, $\hat{g}_{ALGN}=P_{U}g$, is biased, with $\mathbb{E}[\hat{g}_{ALG}]=\frac{q}{d}g$. Here, we compute the MSE directly. Let $\hat{g}_{U}=P_{U}g$.
$\text{MSE}(\hat{g}_{ALG})$ $=\mathbb{E}[\|\hat{g}_{U}-g\|^{2}]=\mathbb{E}[\|(P_{U}-I)g\|^{2}]$
$=\mathbb{E}[g^{T}(P_{U}-I)^{T}(P_{U}-I)g].$
Since $P_{U}$ is a projection matrix, it is symmetric ($P_{U}^{T}=P_{U}$) and idempotent ($P_{U}^{2}=P_{U}$). Therefore, $(P_{U}-I)^{2}=P_{U}^{2}-2P_{U}+I=P_{U}-2P_{U}+I=I-P_{U}$.
$\text{MSE}(\hat{g}_{ALG})$ $=\mathbb{E}[g^{T}(I-P_{U})g]=g^{T}\mathbb{E}[I-P_{U}]g$
$=g^{T}(I-\mathbb{E}[P_{U}])g.$
Using the known result that the expectation of a projection onto a random $q$-dimensional subspace is $\mathbb{E}[P_{U}]=\frac{q}{d}I$, we have:
$\text{MSE}(\hat{g}_{ALIGN})=g^{T}(I-\frac{q}{d}I)g=\left(1-\frac{q}{d}\right)\|g\|^{2}=\frac{d-q}{d}\|g\|^{2}.$
Proof of Theorem 4.1
Proof.
The first part of the proof establishes the one-step contraction factor. From the $L$-smoothness of $f$, we have the descent lemma:
$f(x_{t+1})\leq f(x_{t})+\langle\nabla f(x_{t}),x_{t+1}-x_{t}\rangle+\frac{L}{2}\|x_{t+1}-x_{t}\|^{2}$
Substituting the update rule $x_{t+1}-x_{t}=-\eta_{t}\hat{g}_{AVG}(x_{t})$ and taking the expectation conditional on $x_{t}$:
$\mathbb{E}[f(x_{t+1})|x_{t}]\leq f(x_{t})-\eta_{t}\langle\nabla f(x_{t}),\mathbb{E}[\hat{g}_{AVG}(x_{t})]\rangle+\frac{L\eta_{t}^{2}}{2}\mathbb{E}[\|\hat{g}_{AVG}(x_{t})\|^{2}]$
Using the properties $\mathbb{E}[\hat{g}_{AVG}(x_{t})]=\nabla f(x_{t})$ and $\mathbb{E}[\|\hat{g}_{AVG}(x_{t})\|^{2}]=\frac{q_{t}+d+1}{q_{t}}\|\nabla f(x_{t})\|^{2}$, and setting the optimal step size $\eta_{t}^{*}=\frac{q_{t}}{L(q_{t}+d+1)}$, we obtain the one-step progress:
$\mathbb{E}[f(x_{t+1})|x_{t}]\leq f(x_{t})-\frac{q_{t}}{2L(q_{t}+d+1)}\|\nabla f(x_{t})\|^{2}$
By $\mu$-strong convexity, $\|\nabla f(x_{t})\|^{2}\geq 2\mu(f(x_{t})-f(x_{*}))$. Let $\delta_{t}=E[f(x_{t})-f(x_{*})]$. Taking the full expectation and applying the convexity condition gives:
$\delta_{t+1}\leq\left(1-\frac{\mu q_{t}}{L(q_{t}+d+1)}\right)\delta_{t}$
Unrolling this recurrence from t=T1t=T-1 down to 0:
$\delta_{T}\leq\left(1-\frac{\mu q_{T-1}}{L(q_{T-1}+d+1)}\right)\delta_{T-1}$
$\leq\left(1-\frac{\mu q_{T-1}}{L(q_{T-1}+d+1)}\right)\left(1-\frac{\mu q_{T-2}}{L(q_{T-2}+d+1)}\right)\delta_{T-2}$
$\dots\leq\left(\prod_{t=0}^{T-1}\left(1-\frac{\mu q_{t}}{L(q_{t}+d+1)}\right)\right)\delta_{0}$
This establishes the first part of the theorem.
For the second part, we must find the sequence $\{q_{t}\}_{t=0}^{T-1}$ that minimizes the final error bound under the budget constraint. This is equivalent to minimizing the total contraction factor. Let $\rho_{t}=\frac{\mu q_{t}}{L(q_{t}+d+1)}$. We solve the following optimization problem:
$\min_{\{q_{t}\},T}\prod_{t=0}^{T-1}(1-\rho_{t})\quad\text{s.t.}\sum_{t=0}^{T-1}q_{t}\leq K,\quad 1\leq q_{t}\leq d,\quad q_{t}\in\mathbb{Z}$
Minimizing the product is equivalent to maximizing its logarithm, $\sum_{t=0}^{T-1}\log(1-\rho_{t})$. Since $\log(x)$ is a strictly increasing function, this is equivalent to maximizing $\sum_{t=0}^{T-1}\rho_{t}$. Let $h(q)=\frac{q}{q+d+1}$. The problem becomes:
$\max_{\{q_{t}\},T}\sum_{t=0}^{T-1}\frac{\mu}{L}h(q_{t})\quad\text{s.t.}\sum_{t=0}^{T-1}q_{t}\leq K,\quad 1\leq q_{t}\leq d,\quad q_{t}\in\mathbb{Z}$
The function $h(q)$ is concave for $q>0$, as its second derivative $h^{\prime\prime}(q)=\frac{-2(d+1)}{(q+d+1)^{3}}$ is negative. By Jensen’s inequality, for a fixed number of iterations $T$, the sum $\sum_{t=0}^{T-1}h(q_{t})$ is maximized when the $q_{t}$ values are as equal as possible. Furthermore, because $h(q)/q$ is a decreasing function of $q$, we gain more ”value per query” for smaller $q$. This implies we should maximize the number of terms in the sum, i.e., maximize $T$. The maximum possible $T$ is $K$, which occurs when we choose $q_{t}=1$ for all $t=0,\dots,K-1$. This strategy satisfies both the concavity argument (all $q_{t}$ are equal) and the budget constraint, and is therefore optimal. ∎
Proof of Theorem 4.2
Proof.
As in A.1, we establish the one-step contraction factor. Using the properties $\mathbb{E}[\hat{g}_{ALG}(x_{t})]=\frac{q_{t}}{d}\nabla f(x_{t})$ and $\mathbb{E}[\|\hat{g}_{ALG}(x_{t})\|^{2}]=\frac{q_{t}}{d}\|\nabla f(x_{t})\|^{2}$ with the optimal step size $\eta_{t}^{*}=1/L$, we find:
$\mathbb{E}[f(x_{t+1})|x_{t}]\leq f(x_{t})-\frac{q_{t}}{2Ld}\|\nabla f(x_{t})\|^{2}$
Applying the $\mu$-strong convexity condition and taking the full expectation yields:
$E[f(x_{t+1})-f(x_{*})]\leq\left(1-\frac{\mu q_{t}}{Ld}\right)(E[f(x_{t})]-f(x_{*}))$
Unrolling this recurrence, just as in the proof of Theorem 4.1, gives the first result.
For the second part, we must find the allocation $\{q_{t}\}$ that minimizes the final error bound. This is equivalent to solving the optimization problem:
$\min_{\{q_{t}\},T}\prod_{t=0}^{T-1}\left(1-\frac{\mu q_{t}}{Ld}\right)\quad\text{s.t.}\sum_{t=0}^{T-1}q_{t}\leq K,\quad 1\leq q_{t}\leq d,\quad q_{t}\in\mathbb{Z}$
Let $c=\mu/(Ld)$. The objective function $F(q_{0},\dots,q_{T-1})=\prod_{t=0}^{T-1}(1-cq_{t})$ is a product of linear terms. This function is concave over the feasible region (the simplex defined by the budget constraint). A fundamental result in convex optimization is that the minimum of a concave function over a compact convex set is achieved at an extreme point (a vertex) of the set.
The vertices of the allocation simplex correspond to solutions where the budget is concentrated as much as possible. Given the constraints, the extreme points are strategies that use the largest possible query count, $q_{t}=d$, for some number of iterations, and $q_{t}=0$ for others. To satisfy the budget $\sum q_{t}=K$, the optimal strategy is therefore to perform $T=K/d$ iterations (assuming $K$ is a multiple of $d$) with $q_{t}=d$ for each iteration. This concentrates the entire query budget into the fewest, most powerful steps, thus minimizing the product of contraction factors. ∎
Proof of Theorem 4.3
Proof.
We analyze the evolution of the squared distance to the optimum, $\|x_{t}-x_{*}\|^{2}$.
$\|x_{t+1}-x_{*}\|^{2}=\|x_{t}-x_{*}\|^{2}-2\eta_{t}\hat{g}_{AVG}(x_{t})^{T}(x_{t}-x_{*})+\eta_{t}^{2}\|\hat{g}_{AVG}(x_{t})\|^{2}$
Taking the expectation conditional on $x_{t}$ and using the properties of the ZO-Avg estimator, we get:
$E[\|x_{t+1}-x_{*}\|^{2}|x_{t}]=\|x_{t}-x_{*}\|^{2}-2\eta_{t}\nabla f(x_{t})^{T}(x_{t}-x_{*})+\eta_{t}^{2}\frac{q_{t}+d+1}{q_{t}}\|\nabla f(x_{t})\|^{2}$
By convexity, we have $\nabla f(x_{t})^{T}(x_{t}-x_{*})\geq f(x_{t})-f(x_{*})$. Substituting the step size $\eta_{t}=\frac{q_{t}}{L(q_{t}+d+1)}$ yields:
$E[\|x_{t+1}-x_{*}\|^{2}|x_{t}]\leq\|x_{t}-x_{*}\|^{2}-\frac{2q_{t}}{L(q_{t}+d+1)}(f(x_{t})-f(x_{*}))+\frac{q_{t}}{L^{2}(q_{t}+d+1)}\|\nabla f(x_{t})\|^{2}$
From the one-step analysis in Theorem 4.1, we know that $E[f(x_{t})-f(x_{t+1})|x_{t}]\geq\frac{q_{t}}{2L(q_{t}+d+1)}\|\nabla f(x_{t})\|^{2}$. This implies $\frac{q_{t}}{L(q_{t}+d+1)}\|\nabla f(x_{t})\|^{2}\leq 2(f(x_{t})-E[f(x_{t+1})|x_{t}])$. Substituting this into the inequality gives:
$E[\|x_{t+1}-x_{*}\|^{2}|x_{t}]\leq\|x_{t}-x_{*}\|^{2}-\frac{2q_{t}}{L(q_{t}+d+1)}(f(x_{t})-f(x_{*}))+\frac{1}{2L}(f(x_{t})-E[f(x_{t+1})|x_{t}])$
Rearranging and taking the full expectation, we let $\delta_{t}=E[f(x_{t})-f(x_{*})]$ and $D_{t}=E[\|x_{t}-x_{*}\|^{2}]$:
$\frac{2q_{t}}{L(q_{t}+d+1)}\delta_{t}\leq D_{t}-D_{t+1}+\frac{1}{2L}(\delta_{t}-\delta_{t+1})$
Summing from $t=0$ to T1T-1:
$\sum_{t=0}^{T-1}\frac{2q_{t}}{L(q_{t}+d+1)}\delta_{t}\leq D_{0}-D_{T}+\frac{1}{2L}(\delta_{0}-\delta_{T})\leq\|x_{0}-x_{*}\|^{2}+\frac{1}{2L}(f(x_{0})-f(x_{*}))$
Since $\delta_{t}$ is a non-increasing sequence, $\delta_{t}\geq\delta_{T}$. Therefore:
$\delta_{T}\left(\sum_{t=0}^{T-1}\frac{2q_{t}}{L(q_{t}+d+1)}\right)\leq\|x_{0}-x_{*}\|^{2}+\frac{1}{2L}(f(x_{0})-f(x_{*}))$
Rearranging gives the first result. To obtain the tightest bound on $\delta_{T}$, we must maximize the sum in the denominator, $\sum_{t=0}^{T-1}\frac{q_{t}}{q_{t}+d+1}$, subject to $\sum q_{t}\leq K$. As established in the proof of Theorem 4.1, the function $h(q)=q/(q+d+1)$ is concave, and the value per query, $h(q)/q$, is maximized at $q=1$. Thus, the optimal strategy is to set $T=K$ and $q_{t}=1$ for all $t$. This yields the second result:
$E[f(x_{K})-f(x_{*})]$ $\leq\frac{L(\|x_{0}-x_{*}\|^{2}+\frac{2}{L}(f(x_{0})-f(x_{*})))}{\sum_{t=0}^{K-1}\frac{2}{d+2}}$
$=\frac{L(d+2)(\|x_{0}-x_{*}\|^{2}+\frac{2}{L}(f(x_{0})-f(x_{*})))}{2K}$
This completes the proof. ∎
Proof of Theorem 4.4
Proof.
Following the same initial steps as in the proof of Theorem 4.3, we start with the squared distance:
$E[\|x_{t+1}-x_{*}\|^{2}|x_{t}]=\|x_{t}-x_{*}\|^{2}-2\eta_{t}E[\hat{g}_{ALG}(x_{t})]^{T}(x_{t}-x_{*})+\eta_{t}^{2}E[\|\hat{g}_{ALG}(x_{t})\|^{2}]$
Using the properties of ZO-Align ($E[\hat{g}_{ALG}]=\frac{q_{t}}{d}\nabla f(x_{t})$, $E[\|\hat{g}_{ALG}\|^{2}]=\frac{q_{t}}{d}\|\nabla f(x_{t})\|^{2}$), step size $\eta_{t}=1/L$, and the convexity property $\nabla f(x_{t})^{T}(x_{t}-x_{*})\geq f(x_{t})-f(x_{*})$:
$E[\|x_{t+1}-x_{*}\|^{2}|x_{t}]\leq\|x_{t}-x_{*}\|^{2}-\frac{2q_{t}}{Ld}(f(x_{t})-f(x_{*}))+\frac{q_{t}}{L^{2}d}\|\nabla f(x_{t})\|^{2}$
From the one-step analysis in Theorem 4.2, we have $E[f(x_{t})-f(x_{t+1})|x_{t}]\geq\frac{q_{t}}{2Ld}\|\nabla f(x_{t})\|^{2}$, which implies $\frac{q_{t}}{Ld}\|\nabla f(x_{t})\|^{2}\leq 2(f(x_{t})-E[f(x_{t+1})|x_{t}])$. Substituting this gives:
$E[\|x_{t+1}-x_{*}\|^{2}|x_{t}]\leq\|x_{t}-x_{*}\|^{2}-\frac{2q_{t}}{Ld}(f(x_{t})-f(x_{*}))+\frac{1}{L}(f(x_{t})-E[f(x_{t+1})|x_{t}])$
Rearranging, taking the full expectation, and summing from $t=0$ to T1T-1:
$\sum_{t=0}^{T-1}\frac{2q_{t}}{Ld}E[f(x_{t})-f(x_{*})]\leq\|x_{0}-x_{*}\|^{2}+\frac{1}{L}(f(x_{0})-f(x_{*}))$
By convexity of $f$, the sequence $E[f(x_{t})-f(x_{*})]$ is non-increasing. Thus, we can lower bound the left side using the final error, $E[f(x_{T})-f(x_{*})]$:
$E[f(x_{T})-f(x_{*})]\left(\sum_{t=0}^{T-1}\frac{2q_{t}}{Ld}\right)\leq\|x_{0}-x_{*}\|^{2}+\frac{1}{L}(f(x_{0})-f(x_{*}))$
Rearranging gives the final bound. ∎
Proof of Theorem 4.5
Proof.
From the $L$-smoothness of $f$, we have the one-step progress from the proof of Theorem 4.1:
$E[f(x_{t+1})|x_{t}]\leq f(x_{t})-\frac{q_{t}}{2L(q_{t}+d+1)}\|\nabla f(x_{t})\|^{2}$
Taking the full expectation and rearranging gives:
$\frac{q_{t}}{2L(q_{t}+d+1)}\mathbb{E}[\|\nabla f(x_{t})\|^{2}]\leq\mathbb{E}[f(x_{t})]-\mathbb{E}[f(x_{t+1})]$
Summing from $t=0$ to T1T-1:
$\sum_{t=0}^{T-1}\frac{q_{t}}{2L(q_{t}+d+1)}\mathbb{E}[\|\nabla f(x_{t})\|^{2}]\leq\sum_{t=0}^{T-1}(\mathbb{E}[f(x_{t})]-\mathbb{E}[f(x_{t+1})])$
The right-hand side is a telescoping sum, which simplifies to $\mathbb{E}[f(x_{0})]-\mathbb{E}[f(x_{T})]\leq f(x_{0})-f^{*}$. For the left-hand side, we have:
$\left(\min_{t=0,...,T-1}\mathbb{E}[\|\nabla f(x_{t})\|^{2}]\right)\left(\sum_{t=0}^{T-1}\frac{q_{t}}{2L(q_{t}+d+1)}\right)\leq\sum_{t=0}^{T-1}\frac{q_{t}}{2L(q_{t}+d+1)}\mathbb{E}[\|\nabla f(x_{t})\|^{2}]$
Combining these gives the first result. To optimize the bound under a fixed budget $K$, we must maximize the denominator $\sum_{t=0}^{T-1}\frac{q_{t}}{q_{t}+d+1}$. As established in Theorem 4.1, this sum is maximized by choosing $q_{t}=1$ for $T=K$ iterations. Substituting this into the bound yields the final rate. ∎
Proof of Theorem 4.6
Proof.
The proof follows the same structure as for Theorem 4.5. From the one-step analysis in Theorem 4.2, we have:
$E[f(x_{t+1})|x_{t}]\leq f(x_{t})-\frac{q_{t}}{2Ld}\|\nabla f(x_{t})\|^{2}$
Taking the full expectation, rearranging, and summing from $t=0$ to T1T-1:
$\sum_{t=0}^{T-1}q_{t}E[\|\nabla f(x_{t})\|^{2}]\leq 2Ld\sum_{t=0}^{T-1}(E[f(x_{t})]-E[f(x_{t+1})])\leq 2Ld(f(x_{0})-f^{*})$
Dividing by $\sum q_{t}$ and using the fact that $\min_{t}E[\|\nabla f(x_{t})\|^{2}]\leq\frac{1}{T}\sum E[\|\nabla f(x_{t})\|^{2}]$ gives the result. The rate depends directly on the total number of queries, $\sum q_{t}$, which is maximized by using the full budget $K$. ∎
Proof of Lemma 4.8
Proof.
By the triangle inequality and the property $\|a+b\|^{2}\leq 2\|a\|^{2}+2\|b\|^{2}$, we have:
$\mathbb{E}_{\xi}[\|\nabla F(x,\xi)\|^{2}]\leq 2\mathbb{E}_{\xi}[\|\nabla F(x,\xi)-\nabla F(x^{*},\xi)\|^{2}]+2\mathbb{E}_{\xi}[\|\nabla F(x^{*},\xi)\|^{2}].$
Since each $F(x,\xi)$ is L-smooth, we have $\|\nabla F(x,\xi)-\nabla F(x^{*},\xi)\|\leq L\|x-x^{*}\|$. Also, from the convexity of $f(x)$, we have the standard result $\|x-x^{*}\|^{2}\leq\frac{2}{L}(f(x)-f(x^{*}))$. Combining these gives:
$\mathbb{E}_{\xi}[\|\nabla F(x,\xi)-\nabla F(x^{*},\xi)\|^{2}]\leq L^{2}\|x-x^{*}\|^{2}\leq 2L(f(x)-f(x^{*})).$
The second term is bounded by Assumption 1. Substituting both bounds yields the result:
$\mathbb{E}_{\xi}[\|\nabla F(x,\xi)\|^{2}]\leq 2(2L(f(x)-f(x^{*})))+2\sigma^{2}=4L(f(x)-f(x^{*}))+2\sigma^{2}.$
Proof of Theorem 4.9
Proof.
We start by analyzing the evolution of the squared distance to the optimum, $\|x_{t}-x^{*}\|^{2}$. The update rule is $x_{t+1}=x_{t}-\eta_{t}\hat{g}_{AVG}(x_{t})$.
$\|x_{t+1}-x^{*}\|^{2}=\|x_{t}-x^{*}\|^{2}-2\eta_{t}\hat{g}_{AVG}(x_{t})^{T}(x_{t}-x^{*})+\eta_{t}^{2}\|\hat{g}_{AVG}(x_{t})\|^{2}$
Let $\mathbb{E}_{t}[\cdot]=\mathbb{E}[\cdot|x_{t}]$ denote the expectation over the random direction matrix $U_{t}$ and the stochastic sample $\xi_{t}$. We have $\mathbb{E}_{t}[\hat{g}_{AVG}(x_{t})]=\nabla f(x_{t})$. From Proposition 3.3 and Lemma 4.8, the expected squared norm of the estimator is bounded as:
$\mathbb{E}_{t}[\|\hat{g}_{AVG}(x_{t})\|^{2}]=\frac{q_{t}+d+1}{q_{t}}\mathbb{E}_{\xi_{t}}[\|\nabla F(x_{t},\xi_{t})\|^{2}]\leq\frac{q_{t}+d+1}{q_{t}}(4L(f(x_{t})-f(x^{*}))+2\sigma^{2})$
Taking the conditional expectation of the distance expansion and using the convexity of $f$, i.e., $f(x_{t})-f(x^{*})\leq\nabla f(x_{t})^{T}(x_{t}-x^{*})$:
$\mathbb{E}_{t}[\|x_{t+1}-x^{*}\|^{2}]$ $\leq\|x_{t}-x^{*}\|^{2}-2\eta_{t}\nabla f(x_{t})^{T}(x_{t}-x^{*})+\eta_{t}^{2}\mathbb{E}_{t}[\|\hat{g}_{AVG}(x_{t})\|^{2}]$
$\leq\|x_{t}-x^{*}\|^{2}-2\eta_{t}(f(x_{t})-f(x^{*}))+\eta_{t}^{2}\frac{q_{t}+d+1}{q_{t}}(4L(f(x_{t})-f(x^{*}))+2\sigma^{2})$
$=\|x_{t}-x^{*}\|^{2}+\left(4L\eta_{t}^{2}\frac{q_{t}+d+1}{q_{t}}-2\eta_{t}\right)(f(x_{t})-f(x^{*}))+2\eta_{t}^{2}\frac{q_{t}+d+1}{q_{t}}\sigma^{2}$
Given the step-size condition $\eta_{t}\leq\frac{1}{4L}$, we have $4L\eta_{t}\leq 1$. We can bound the coefficient of the suboptimality gap:
$4L\eta_{t}^{2}\frac{q_{t}+d+1}{q_{t}}-2\eta_{t}=\frac{\eta_{t}}{q_{t}}(4L\eta_{t}(q_{t}+d+1)-2q_{t})\leq\frac{\eta_{t}}{q_{t}}(1\cdot(q_{t}+d+1)-2q_{t})=\frac{\eta_{t}(d+1-q_{t})}{q_{t}}$
Substituting this bound back into the inequality for the distance:
$\mathbb{E}_{t}[\|x_{t+1}-x^{*}\|^{2}]\leq\|x_{t}-x^{*}\|^{2}+\frac{\eta_{t}(d+1-q_{t})}{q_{t}}(f(x_{t})-f(x^{*}))+2\eta_{t}^{2}\frac{q_{t}+d+1}{q_{t}}\sigma^{2}$
Rearranging, taking the full expectation with $\delta_{t}=\mathbb{E}[f(x_{t})-f(x^{*})]$ and $D_{t}=\mathbb{E}[\|x_{t}-x^{*}\|^{2}]$:
$\frac{\eta_{t}(d+1-q_{t})}{q_{t}}\delta_{t}\leq D_{t}-D_{t+1}+2\eta_{t}^{2}\frac{q_{t}+d+1}{q_{t}}\sigma^{2}$
Summing this inequality from $t=0$ to T1T-1:
$\sum_{t=0}^{T-1}\frac{\eta_{t}(d+1-q_{t})}{q_{t}}\delta_{t}$ $\leq\sum_{t=0}^{T-1}(D_{t}-D_{t+1})+\sum_{t=0}^{T-1}2\eta_{t}^{2}\frac{q_{t}+d+1}{q_{t}}\sigma^{2}$
$\leq D_{0}-D_{T}+2\sigma^{2}\sum_{t=0}^{T-1}\eta_{t}^{2}\frac{q_{t}+d+1}{q_{t}}\leq\|x_{0}-x^{*}\|^{2}+2\sigma^{2}\sum_{t=0}^{T-1}\eta_{t}^{2}\frac{q_{t}+d+1}{q_{t}}$
Let $\alpha_{t}=\frac{\eta_{t}(d+1-q_{t})}{q_{t}}$. By convexity of $f$ and Jensen’s inequality, the suboptimality of the weighted average iterate $\bar{x}_{T}=\frac{1}{\sum\alpha_{t}}\sum\alpha_{t}x_{t}$ is bounded by $\mathbb{E}[f(\bar{x}_{T})-f(x^{*})]\leq\frac{\sum\alpha_{t}\delta_{t}}{\sum\alpha_{t}}$. This gives:
$\mathbb{E}[f(\bar{x}_{T})-f(x^{*})]\leq\frac{\|x_{0}-x^{*}\|^{2}+2\sigma^{2}\sum_{t=0}^{T-1}\eta_{t}^{2}\frac{q_{t}+d+1}{q_{t}}}{\sum_{t=0}^{T-1}\eta_{t}\frac{d+1-q_{t}}{q_{t}}}$
Substituting the step size $\eta_{t}=\eta_{0}/\sqrt{t+1}$ yields the bound stated in the theorem.
To find the optimal query allocation, we analyze this upper bound, $U(\{q_{t}\})$, under the budget constraint $\sum_{t=0}^{T-1}q_{t}\leq K$. We consider a constant query count, $q_{t}=q$, which implies $T=K/q$. For large $K$, we use the asymptotic approximations $\sum_{t=0}^{T-1}\frac{1}{t+1}\approx\ln(T)$ and $\sum_{t=0}^{T-1}\frac{1}{\sqrt{t+1}}\approx 2\sqrt{T}$. Substituting $T=K/q$:
$U_{asym}(q)\approx\frac{\|x_{0}-x^{*}\|^{2}+2\eta_{0}^{2}\sigma^{2}\left(\frac{q+d+1}{q}\right)\ln(K/q)}{2\eta_{0}\left(\frac{d+1-q}{q}\right)\sqrt{K/q}}$
Rearranging terms to analyze the dependence on $q\in[1,d]$:
$U_{asym}(q)\approx\frac{1}{2\eta_{0}\sqrt{K}}\left[\frac{q^{3/2}}{d+1-q}\|x_{0}-x^{*}\|^{2}+2\eta_{0}^{2}\sigma^{2}\frac{\sqrt{q}(q+d+1)}{d+1-q}\ln\left(\frac{K}{q}\right)\right]$
To find the minimum, we analyze the monotonicity of the functions of $q$. The first term is proportional to $g_{1}(q)=\frac{q^{3/2}}{d+1-q}$. Its derivative is $g_{1}^{\prime}(q)=\frac{q^{1/2}(\frac{3}{2}(d+1)-\frac{1}{2}q)}{(d+1-q)^{2}}$, which is positive for $q\in[1,d]$. Thus, the first term is monotonically increasing in $q$. The second term is proportional to $g_{2}(q)=\frac{\sqrt{q}(q+d+1)}{d+1-q}(\ln K-\ln q)$. For a sufficiently large budget $K$, this term is also monotonically increasing in $q$. Since both terms in the asymptotic upper bound are monotonically increasing functions of $q$, the total bound $U(q)$ is minimized at the smallest possible integer value for the query count, which is $q=1$. This strategy implies using the full budget for the number of iterations, $T=K$. ∎
Proof of Theorem 4.10
Proof.
We start by analyzing the evolution of the squared distance to the optimum, $\|x_{t}-x^{*}\|^{2}$. The update rule is $x_{t+1}=x_{t}-\eta_{t}\hat{g}_{ALG}(x_{t})$.
$\|x_{t+1}-x^{*}\|^{2}=\|x_{t}-x^{*}\|^{2}-2\eta_{t}\hat{g}_{ALG}(x_{t})^{T}(x_{t}-x^{*})+\eta_{t}^{2}\|\hat{g}_{ALG}(x_{t})\|^{2}$
Let $\mathbb{E}_{t}[\cdot]=\mathbb{E}[\cdot|x_{t}]$ denote the expectation over the random direction matrix $U_{t}$ and the stochastic sample $\xi_{t}$. Using the properties of the ZO-Align estimator, we have $\mathbb{E}_{t}[\hat{g}_{ALG}(x_{t})]=\frac{q_{t}}{d}\nabla f(x_{t})$ and $\mathbb{E}_{t}[\|\hat{g}_{ALG}(x_{t})\|^{2}]=\frac{q_{t}}{d}\mathbb{E}_{\xi_{t}}[\|\nabla F(x_{t},\xi_{t})\|^{2}]$. Using Lemma 4.8, this becomes:
$\mathbb{E}_{t}[\|\hat{g}_{ALG}(x_{t})\|^{2}]\leq\frac{q_{t}}{d}(4L(f(x_{t})-f(x^{*}))+2\sigma^{2})$
Taking the conditional expectation of the distance expansion and using the convexity of $f$, i.e., $f(x_{t})-f(x^{*})\leq\nabla f(x_{t})^{T}(x_{t}-x^{*})$:
$\mathbb{E}_{t}[\|x_{t+1}-x^{*}\|^{2}]$ $\leq\|x_{t}-x^{*}\|^{2}-2\eta_{t}\frac{q_{t}}{d}\nabla f(x_{t})^{T}(x_{t}-x^{*})+\eta_{t}^{2}\mathbb{E}_{t}[\|\hat{g}_{ALG}(x_{t})\|^{2}]$
$\leq\|x_{t}-x^{*}\|^{2}-\frac{2\eta_{t}q_{t}}{d}(f(x_{t})-f(x^{*}))+\eta_{t}^{2}\frac{q_{t}}{d}(4L(f(x_{t})-f(x^{*}))+2\sigma^{2})$
$=\|x_{t}-x^{*}\|^{2}+\left(\frac{4L\eta_{t}^{2}q_{t}}{d}-\frac{2\eta_{t}q_{t}}{d}\right)(f(x_{t})-f(x^{*}))+\frac{2\eta_{t}^{2}q_{t}}{d}\sigma^{2}$
Given the step-size condition $\eta_{t}\leq\frac{1}{4L}$, we have $4L\eta_{t}\leq 1$. We can bound the coefficient of the suboptimality gap:
$\frac{4L\eta_{t}^{2}q_{t}}{d}-\frac{2\eta_{t}q_{t}}{d}=\frac{\eta_{t}q_{t}}{d}(4L\eta_{t}-2)\leq\frac{\eta_{t}q_{t}}{d}(1-2)=-\frac{\eta_{t}q_{t}}{d}$
Substituting this bound back into the inequality for the distance:
$\mathbb{E}_{t}[\|x_{t+1}-x^{*}\|^{2}]\leq\|x_{t}-x^{*}\|^{2}-\frac{\eta_{t}q_{t}}{d}(f(x_{t})-f(x^{*}))+\frac{2\eta_{t}^{2}q_{t}}{d}\sigma^{2}$
Rearranging, taking the full expectation with $\delta_{t}=\mathbb{E}[f(x_{t})-f(x^{*})]$ and $D_{t}=\mathbb{E}[\|x_{t}-x^{*}\|^{2}]$:
$\frac{\eta_{t}q_{t}}{d}\delta_{t}\leq D_{t}-D_{t+1}+\frac{2\eta_{t}^{2}q_{t}}{d}\sigma^{2}$
Summing this inequality from $t=0$ to T1T-1:
$\sum_{t=0}^{T-1}\frac{\eta_{t}q_{t}}{d}\delta_{t}$ $\leq\sum_{t=0}^{T-1}(D_{t}-D_{t+1})+\sum_{t=0}^{T-1}\frac{2\eta_{t}^{2}q_{t}}{d}\sigma^{2}$
$\leq D_{0}-D_{T}+\frac{2\sigma^{2}}{d}\sum_{t=0}^{T-1}\eta_{t}^{2}q_{t}\leq\|x_{0}-x^{*}\|^{2}+\frac{2\sigma^{2}}{d}\sum_{t=0}^{T-1}\eta_{t}^{2}q_{t}$
Let $\alpha_{t}=\frac{\eta_{t}q_{t}}{d}$. By convexity of $f$ and Jensen’s inequality, the suboptimality of the weighted average iterate $\bar{x}_{T}=\frac{1}{\sum\alpha_{t}}\sum\alpha_{t}x_{t}$ is bounded by $\mathbb{E}[f(\bar{x}_{T})-f(x^{*})]\leq\frac{\sum\alpha_{t}\delta_{t}}{\sum\alpha_{t}}$. This gives:
$\mathbb{E}[f(\bar{x}_{T})-f(x^{*})]\leq\frac{\|x_{0}-x^{*}\|^{2}+\frac{2\sigma^{2}}{d}\sum_{t=0}^{T-1}\eta_{t}^{2}q_{t}}{\sum_{t=0}^{T-1}\frac{\eta_{t}q_{t}}{d}}=\frac{d\|x_{0}-x^{*}\|^{2}+2\sigma^{2}\sum_{t=0}^{T-1}\eta_{t}^{2}q_{t}}{\sum_{t=0}^{T-1}\eta_{t}q_{t}}$
Substituting the step size $\eta_{t}=\eta_{0}/\sqrt{t+1}$ yields the bound stated in the theorem.
To optimize the bound, we consider the case where the first term in the numerator (the ”bias” term related to the initial distance) dominates the second term (the ”variance” term related to $\sigma^{2}$). This is common in the initial phase of optimization. The goal is to minimize the bound by maximizing the denominator, $\sum_{t=0}^{T-1}\frac{\eta_{0}q_{t}}{\sqrt{t+1}}$, subject to the total budget constraint $\sum_{t=0}^{T-1}q_{t}\leq K$. Let’s analyze the objective $\max\sum_{t=0}^{T-1}\frac{q_{t}}{\sqrt{t+1}}$. The term $\frac{1}{\sqrt{t+1}}$ represents the ”value” or ”effectiveness” of queries used at iteration $t$. Since this value is a strictly decreasing function of $t$, we should allocate our query budget to the earliest possible iterations to achieve the maximum sum. This implies a greedy strategy: use the maximum allowed queries, $q_{t}=d$, for the first $T=K/d$ iterations. This concentrates the entire budget $K$ where it is most effective, thus maximizing the denominator and providing the tightest convergence bound. ∎
A.2 Experiment Set Up
Objective Functions
For the strongly convex quadratic function, $f(x)=\frac{1}{2}x^{T}Ax+b^{T}x$, the positive definite matrix $A$ was generated by first creating a random matrix $M$ with entries sampled from a standard normal distribution. $A$ was then computed as $A=M^{T}M+\epsilon I$, where $\epsilon$ is a small positive constant to ensure the matrix is well-conditioned. The vector $b$ was also sampled from a standard normal distribution.
For the convex logistic regression task, synthetic data was generated by sampling feature vectors $a_{i}\in\mathbb{R}^{1000}$ from a standard normal distribution. A ground-truth weight vector $w_{true}$ was randomly created, and the binary labels $y_{i}\in\{-1,1\}$ were then generated based on the sign of $a_{i}^{T}w_{true}$.
The non-convex objective was the standard Rosenbrock function.
The stochastic convex objective used a logistic regression loss, where mini-batches were sampled at each iteration to compute the function value.
Implementation and Evaluation
We evaluated the ZO-Avg and ZO-Align estimators using various query block sizes $q$. Step sizes $\eta_{t}$ were set according to the optimal values derived in our theory: for ZO-Avg, $\eta_{t}=\frac{q_{t}}{L(q_{t}+d+1)}$, and for ZO-Align, $\eta_{t}=1/L$. For the stochastic setting, a standard diminishing step-size rule was applied. Each algorithm was tested under two total query budgets: a large budget of K=20,000 and a constrained budget of K=500. All reported results are an average of 10 independent runs to ensure reliability.
Hardware and Software
Our experiments were conducted using Python 3.9 and PyTorch 1.12. The underlying system was Ubuntu 18.04, equipped with an Intel Xeon Gold 5320 CPU and two NVIDIA RTX 3090 GPUs.